OpenStudy (anonymous):

solve ln 5+ln (2x)=5 round your answer to the nearest hundredth. Please help.

Nnesha (nnesha):

write in single ln form product rule \[\large \rm ln x + \ln y = \ln ( x \times y )\]

OpenStudy (anonymous):

Could you go on? because I'm still confused.

Nnesha (nnesha):

let's say x represent 5 and 2x represent y how would you condense ln in single ln form ? just like in this equation \[\large \rm ln x + \ln y = \ln ( x \times y )\]

OpenStudy (anonymous):

I keep getting 1/2

Nnesha (nnesha):

\[\large \rm ln x + \ln y = \ln ( x \times y )\] \[\large\rm ln 5 + \ln (2x) = ?\]

Nnesha (nnesha):

show your work how did you get 1/2 ?

OpenStudy (anonymous):

Well it goes like ln(2x)=−5+5 ln(2x)=0 (e)0=2x e0=2x 1=2x 2x=1 x=1/2

Nnesha (nnesha):

alright first step is wrong why did you subtract 5 ?

Nnesha (nnesha):

it's ln 5 at left side not just 5

OpenStudy (anonymous):

My choices are 1.1484 2.11.17 3.71.71 4.0.02 also since Since 5 does not contain the variable to solve for, we move it to the right-hand side of the equation by subtracting 5 from both sides. If I'm wrong can you direct me in the direction of the right answer by showing me how to do it?

Nnesha (nnesha):

alright nope you can't subtract 5 bec it's ln(5)

OpenStudy (anonymous):

Okay, well can you show me, please ?

Nnesha (nnesha):

first of all apply product property \[\large \rm ln x + \ln y = \ln ( x \times y )\] \[\large\rm ln 5 + \ln (2x) = ln(5 \times2x) \] there is a plus sign so you have to apply product property

Nnesha (nnesha):

\[\huge\rm ln (5 \times 2x) =5\] now take e both side do it let me know what yo uget

OpenStudy (anonymous):

3.85244

OpenStudy (anonymous):

Wait I did something wrong

Nnesha (nnesha):

yes try again :-)

OpenStudy (anonymous):

Sorry, I'm not sure what to do. I'm pretty brain dead because I have been working on my math final since like 7 am.

Nnesha (nnesha):

math final ?

Nnesha (nnesha):

is it from a test ?? :-)

OpenStudy (anonymous):

Mhm, I have about 16 questions left out of 50.

Nnesha (nnesha):

sorry...

OpenStudy (kidrah69):

We help study for tests we DO NOT help you ON the test. Please go find a diffrent website to cheat from because we will not help you with a test refer the the Terms and conditions which YOU agreed to before you decided to come to this site. thanks.

eclipsedstar (eclipsedstar):

http://goo.gl/NJQwpm

TheSmartOne (thesmartone):

http://www.nig.gr/brJ

OpenStudy (anonymous):

I didn't feel like it was cheating, I was just asking for explanation rather than the answer, you know?

OpenStudy (anonymous):

Well I'll close the question, I wasn't aware, sorry.

OpenStudy (kidrah69):

Would you teacher want to see what U know or what we know?